# AI-Empowered Integrated Sensing and Communications

### v1

@@ Abstract
Integrating sensing and communication (ISAC) can help overcome the challenges of limited spectrum and expensive hardware, leading to improved energy and cost efficiency. While full cooperation between sensing and communication can result in significant performance gains, achieving optimal performance requires efficient designs of unified waveforms and beamformers for joint sensing and communication. Sophisticated statistical signal processing and multi-objective optimization techniques are necessary to balance the competing design requirements of joint sensing and communication tasks. Since model-based analytical approaches may be suboptimal or overly complex, deep learning emerges as a powerful tool for developing data-driven signal processing algorithms, particularly when optimal algorithms are unknown or when known algorithms are too complex for real-time implementation. Unified waveform and beamformer design problems for ISAC fall into this category, where fundamental design trade-offs exist between sensing and communication performance metrics, and the underlying models may be inadequate or incomplete. This article explores the application of artificial intelligence (AI) in ISAC designs to enhance efficiency and reduce complexity. We emphasize the integration benefits through AI-driven ISAC designs, prioritizing the development of unified waveforms, constellations, and beamforming strategies for both sensing and communication. To illustrate the practical potential of AI-driven ISAC, we present two case studies on waveform and beamforming design, demonstrating how unsupervised learning and neural network–based optimization can effectively balance performance, complexity, and implementation constraints.
Index Terms:
Integrating sensing and communications (ISAC), AI, beamforming, waveform, constellation, radar, sensing, performance metrics, unsupervised learning, autoencoder, 6G.
@@ I Introduction
The performance of sensing tasks—such as estimation, detection, and recognition—is typically quantified via mean squared error (MSE), Cramér-Rao lower bound (CRLB), and receiver operating characteristic (ROC) [3, 4, 5]. On the other hand, the performance of wireless communications is mainly characterized by spectral and energy efficiency and reliability metrics. Consequently, there are fundamental trade-offs between the performance metrics of sensing and communications. In particular, in a Gaussian channel, the main metric for communication data rates—the mutual information (MI)—is related to the minimum mean square error (MMSE) of sensing by $\frac{\partial}{\partial\gamma}I(\gamma)=0.5\mathrm{MMSE}(\gamma),$ where $\gamma$ represents the signal-to-noise ratio (SNR)[6]. Thus, for a Gaussian channel, Gaussian inputs provide an upper bound on both the mutual information and MMSE. In other words, the Gaussian input is the most favorable for communication (as it maximizes $I(\gamma)$), yet the least favorable for sensing (as it maximizes $\mathrm{MMSE}(\gamma)$). This is demonstrated in Fig. 1, where the $\rm MMSE$ and $\rm MI$ are also plotted for two practical constellations, namely binary phase-shift keying (BPSK) and quadrature phase-shift keying (QPSK). It is clear that QPSK is preferred if mutual information is the metric, whereas BPSK is preferred when MMSE is a concern.
@@
Figure 1: Mutual information ($I$ in nats per channel-use) and MMSE versus SNR for an AWGN channel with three different input signals: Gaussian, QPSK, and BPSK. While Gaussian signaling achieves the maximum information rate, it also results in the highest (worst) MMSE.
@@
Given the fundamental trade-offs and the inherent multi-objective optimization requirements, advanced statistical signal processing and multi-objective optimization techniques are necessary to balance the competing design objectives of joint sensing and communication. Significant research efforts have been dedicated to achieving the goals of ISAC through classical model-based designs and theoretical analyses, leveraging communication and statistical signal processing theories [7, 1, 3, 8, 9, 10]. However, designing optimal unified waveforms and beamformers using model-based analytical approaches remains often prohibitively complex and challenging.
@@
Machine learning (ML), and its subset deep learning (DL), has revolutionized wireless physical layer designs [11]. Deep learning is a powerful tool for extracting channel features and leveraging them in various tasks through universal model approximations and solving non-linear multi-objective optimization problems. The use of DL/ML offers significant advantages in this context by enabling performance optimization without relying on explicit models. Instead, DL approaches leverage implicit learning to extract channel parameters directly from high-dimensional data, making them well-suited for the dynamic and unpredictable conditions inherent to ISAC systems.
I-A Motivation
Driven by the limitations of model-based analytical approaches in designing efficient unified beamforming and waveform solutions, this tutorial aims to establish the theoretical and algorithmic foundations of artificial intelligence (AI)-assisted waveform and beamformer designs. These AI-driven approaches can unify and optimize ISAC functionality while enhancing ISAC integration gains. A critical question is when and why learning-based approaches are needed. We answer this in the following.
I-A1 Imperfect Modeling of Mismatches
@@
In the context of ISAC, learning-based approaches become essential when traditional model-based methods fail to account for imperfections and mismatches in models or hardware. These mismatches can stem from imperfect (e.g., estimated or quantized) channel state information (CSI), dynamic environments, interference, hardware limitations, or the inherent complexity of integrating sensing and communication. Learning-based techniques, such as deep learning or reinforcement learning, can adapt to these uncertainties by leveraging real-time data, enabling robust waveform and beamformer designs. This adaptability is crucial for optimizing ISAC performance in scenarios where precise analytical models are either insufficient or computationally intractable.
I-A2 End-to-End Learning
Deep autoencoder-based end-to-end learning [12] offers a powerful approach to system design, as it enables the joint optimization of multiple components without relying on explicit models, making it highly adaptable to the dynamic and uncertain environments characteristic of ISAC. The broad impact and benefits of DL, and autoencoders in particular, are mainly due to their effectiveness in processing high-dimensional data with embedded low-dimensional structures [13]. This characteristic makes them particularly attractive for ISAC problems such as CSI estimation, interference management, constellation design, and waveform design.
For example, from a wireless communications perspective, received pilot signals represent high-dimensional observations typically acquired through sparse mmWave and sub-terahertz (THz) channels, which inherently possess low-dimensional structure. However, the number of such observations is limited by the channel coherence interval, which is constrained by delay and Doppler spreads. Besides, when hybrid beamforming is employed, the baseband output becomes a nonlinear, low-dimensional transformation of these high-dimensional observations, making accurate estimation of radar and communication channel parameters highly challenging.
I-A3 6G Networks Are Underpinned by AI
6G networks are expected to be AI-driven. Embedded AI into networking equipment enables the network to learn, self-manage, operate autonomously, and reduce costs. AI-driven 6G networks leverage machine learning at every layer—from physical layer optimization to network management—facilitating seamless integration of sensing and communication functionalities. ISAC is anticipated to play a pivotal role in 6G systems [3, 14, 4], advanced vehicle-to-everything systems [15, 16, 17], smart industrial Internet of things (IoT) networks [18, 19, 20, 21], and next-generation wireless local area networks [22, 23]. Applications of ISAC include, but are not limited to, localization, tracking, autonomous vehicles, area imaging, human activity detection, environmental monitoring, and surveillance [3, 24, 10]. Incorporating AI into ISAC further enhances the development of highly adaptive, intelligent, and efficient communication systems.
I-B Scope and Related Tutorials
@@
This work presents a tutorial dedicated to AI-based ISAC, with a primary focus on how AI can enhance ISAC design through the joint optimization of waveform and beamforming. We discuss various ML/DL–based optimization approaches for this joint design, highlighting their advantages over traditional analytical and numerical methods used for multi-objective optimization.
Several papers have surveyed the past research contributions to the advancement of ISAC technology [24, 25, 26, 27, 1, 28, 29, 30, 31, 32, 33, 34]. However, none of those tutorials and surveys are dedicated to AI in ISAC. In [1], a survey on perceptive mobile networks explores ISAC functionalities beyond localization, tracking, and object recognition. Such networks serve as radio sensing platforms, enabling smart sensing applications while ensuring uncompromised communication. In [16, 24], the fundamental performance limits of ISAC have been studied. In particular, [24] categorizes existing techniques as device-based ISAC, device-free ISAC, device-free sensing, and device-based sensing. A survey on orthogonal frequency-division multiplexing (OFDM) and orthogonal time frequency space (OTFS) modulation for ISAC is presented in [25].A more comprehensive survey on ISAC waveform and multiple access design can be found in [33]. In [26], an overview of network sensing and sensing-aided communication techniques for cellular systems is provided. Fundamental limits of ISAC are presented in [34].
The integration gains, system and signal design, network resource management for integrated sensing, communication, and computation are discussed in [27]. A survey of interference management techniques for ISAC is presented in [28]. Multi-functional reconfigurable intelligent surface-aided ISAC systems are reviewed in [29]. In [30], intelligent multi-modal sensing-communication integration is surveyed using the concept of machine synesthesia for dynamic environments. A survey on secure ISAC for intelligent connectivity in 6G networks is presented in [31], emphasizing the need for dynamic adaptability and self-learning capabilities. Finally, [32] provides a recent tutorial on sensing with communication signals.
@@ I-C Contribution
We begin with a primer on ISAC, introducing key concepts in both communication and sensing, along with performance metrics relevant to each domain. Radar technology is presented as a a core sensing component, and its integration with communication systems is discussed. We then frame ISAC problems as multi-objective optimization tasks, comparing conventional approaches—such as weighted sum, $\epsilon$-constraint, and Pareto optimization—with modern learning-based solutions, including multi-task learning and multi-objective learning.
A major contribution of this tutorial is its detailed examination of waveform and beamforming design for ISAC and the role AI plays in enhancing it. We cover traditional beamforming techniques in communication, radar, and ISAC and explore learning-based approaches, including supervised, unsupervised, and reinforcement learning, as well as model-based learning techniques like plug-and-play methods and algorithm unrolling.
To make AI-driven ISAC designs more tangible, we present two concrete case studies on AI-empowered waveform and beamforming design.
@@
  • The second case study presents a hybrid beamforming method using a learned optimizer, which optimizes analog and digital precoders by reformulating the problem as constrained least-squares optimization, and solves it via neural network-based projection. Numerical results show that the proposed unrolled hybrid beamforming method achieves spectral efficiency close to optimal digital beamforming, particularly with more receiving antennas.
These studies underscore AI’s potential for efficient and scalable waveform design and beamforming in next-generation wireless systems.
Finally, we outline future research directions and open problems, such as the development of end-to-end ISAC systems, the application of transformer neural networks for unified waveform and beamformer design, and digital twin-aided ISAC architectures. By bridging the gap between traditional signal processing and modern AI-driven techniques, this tutorial serves as a valuable resource for researchers and practitioners at the forefront of ISAC technology.
TABLE I: List of Key Abbreviations.
@@
PMCW Phase modulated continuous wave

### v2

# A Tutorial on AI-Empowered Integrated Sensing and Communications

@@ Abstract
Integrating sensing and communication (ISAC) can help overcome the challenges of limited spectrum and expensive hardware, leading to improved energy and cost efficiency. While full cooperation between sensing and communication can result in significant performance gains, achieving optimal performance requires efficient designs of unified waveforms and beamformers for joint sensing and communication. Sophisticated statistical signal processing and multi-objective optimization techniques are necessary to balance the competing design requirements of joint sensing and communication tasks. As model-based approaches can be suboptimal or too complex, deep learning offers a powerful data-driven alternative, especially when optimal algorithms are unknown or impractical for real-time use. Unified waveform and beamformer design problems for ISAC fall into this category, where fundamental design trade-offs exist between sensing and communication performance metrics, and the underlying models may be inadequate or incomplete. This tutorial paper explores the application of artificial intelligence (AI) to enhance efficiency or reduce complexity in ISAC designs. We emphasize the integration benefits through AI-driven ISAC designs, prioritizing the development of unified waveforms, constellations, and beamforming strategies for both sensing and communication. To illustrate the practical potential of AI-driven ISAC, we present three case studies on waveform, beamforming, and constellation design, demonstrating how unsupervised learning and neural network–based optimization can effectively balance performance, complexity, and implementation constraints.
@@ I Introduction
The performance of sensing tasks, such as estimation, detection, and tracking, is typically quantified via mean squared error (MSE), Cramér-Rao lower bound (CRLB), and receiver operating characteristic [3, 4, 5]. On the other hand, the performance of wireless communications is mainly characterized by spectral and energy efficiency and reliability metrics. There are fundamental trade-offs between the performance metrics of sensing and communications. In particular, in a Gaussian channel, the mutual information (MI), the main metric for the transmission data rates, is related to the minimum mean square error (MMSE) of sensing by $\frac{\partial}{\partial\gamma}I(\gamma)=0.5\mathrm{MMSE}(\gamma),$ where $\gamma$ represents the signal-to-noise ratio (SNR)[6]. Thus, for a Gaussian channel, Gaussian inputs provide an upper bound on both the mutual information and MMSE. In other words, the Gaussian input is the most favorable for communication (as it maximizes $I(\gamma)$), yet the least favorable for sensing (as it maximizes $\mathrm{MMSE}(\gamma)$). This is demonstrated in Fig. 1, where the $\rm MMSE$ and $\rm MI$ are also plotted for two practical constellations, namely binary phase-shift keying (BPSK) and quadrature phase-shift keying (QPSK). It is clear that QPSK is preferred if mutual information is the metric, whereas BPSK is preferred when MMSE is a concern. [^1] In Section VI-C, we illustrate how novel constellations can be designed for ISAC to effectively balance the trade-off between sensing and communication.
@@
Figure 1: Mutual information (II in nats/channel-use) and MMSE versus SNR for an additive white Gaussian noise (AWGN) channel with three different input signals: Gaussian, QPSK, and BPSK. While Gaussian signaling achieves the maximum information rate, it also results in the highest (worst) MMSE.
@@
Given the fundamental trade-offs and the inherent multi-objective optimization requirements, advanced statistical signal processing and multi-objective optimization techniques are necessary to balance the competing design objectives of joint sensing and communication. Significant research efforts have been dedicated to achieving the goals of ISAC through classical model-based designs and theoretical analyses, leveraging communication and statistical signal processing theories [7, 1, 3, 8, 9, 10]. However, designing unified constellations, waveforms and beamformers using model-based analytical approaches remains often prohibitively complex and challenging.
@@
Artificial intelligence (AI) and its subsets, machine learning (ML) and deep learning, have revolutionized wireless physical-layer design. [11]. Particularly, deep learning (DL) is a powerful tool for extracting channel features and leveraging them in various tasks through universal model approximations and solving non-linear multi-objective optimization problems. The use of DL/ML offers significant advantages in this context by enabling performance optimization without relying on explicit models. Instead, DL approaches leverage implicit learning to extract channel parameters directly from high-dimensional data, making them well-suited for the dynamic and unpredictable conditions inherent to sensing and communication systems. DL designs based on data-driven, model-driven, and their hybrids have been investigated for a variety of physical-layer communication and sensing problems [12, 13, 14, 15, 16, 17].
I-A Motivations of AI-Based ISAC
Classical model-based analytical approaches may be suboptimal due to model inadequacy or incompleteness, or become overly complex for ISAC problems because of the fundamental trade-offs between sensing and communication performance metrics, where optimal designs require sophisticated multi-objective optimization. Motivated by these limitations in developing efficient unified beamforming and waveform solutions, this tutorial focuses o providing an overview of efficient AI- empowered ISAC designs, introducing underlying learning- based design foundations, presenting practical case studies, evaluating their performance, and comparing them against the classical model-based analytical counterparts.
A critical question is when and why learning-based approaches are needed in the context of ISAC. Performance gains of DL-based designs over model-based analytical counterparts in physical-layer are prevalent in cases where (i) the optimal algorithms are unknown, (ii) the known algorithms can acquire the optimal solution but are prohibitively complexity for real-time implementation, and (iii) the system models are either inadequate or incomplete [18]. Unified waveform and beamformer design problems for ISAC systems fall into one or more of the aforementioned categories. We elaborate on this in what follows.
I-A1 Model Inadequacy & Hardware Impairments
@@
Learning-based approaches become essential when traditional model-based methods fail to adequately account for imperfections and mismatches in models or hardware. These mismatches can stem from imperfect (e.g., estimated or quantized) channel state information (CSI), dynamic environments, where CSI changes rapidly, hardware limitations, or the inherent complexity of ISAC. Learning-based techniques can adapt to these uncertainties by leveraging real-time data, enabling robust waveform and beamformer designs. This adaptability is crucial for optimizing ISAC performance in scenarios where precise analytical models are either insufficient or computationally intractable. For example, [19] shows that, compared to convex optimization methods, deep learning can significantly enhance the performance of both sensing and communication while reducing complexity. Similarly, [20] demonstrates that learning-based approaches in ISAC outperform traditional optimization techniques, particularly in terms of error rates under hardware impairments. Additional examples can be found in [21].
TABLE I: Existing ISAC surveys and tutorials and their primary focus
[…]
2026 This paper AI-empowered ISAC waveforms
[…]
2025 [23] Sensing, communication, and computation
[…]
2025 [25] Interference management for ISAC
[…]
2023 [27] ISAC signal design and optimization
[…]
2022 [3] ISAC background and key applications
[…]
2022 [1] Integrating sensing to mobile networks
I-A2 Simplifying System Design
DL architectures, such as end-to-end learning [30], offer a powerful alternative to the block-by-block system design. End-to-end learning enables to jointly optimize multiple blocks, e.g., encoding, modulation, and beamforming, thereby revolutionizing and simplifying system design compared to conventional methods. This approach is a significant step toward AI-native air interface which is a long-term goal in AI-based communication system design [31]. The benefits of DL, and autoencoders in particular, stem from their ability to learn low-dimensional structures embedded within high-dimensional data [32]. In ISAC systems, received radar echoes or pilot signals of the communication channel can be viewed as high-dimensional observations of sparse structures, making autoencoders suitable for joint optimization tasks such as channel estimation, target detection, beamforming, and constellation or waveform design in ISAC systems [20]. DL models can also jointly optimize tasks such as range-Doppler estimation, target classification, beam pattern design, and waveform adaptation [21]. Further, when optimization algorithms are too complex or computationally intensive, DL methods such as plug-and-play [33] and algorithm unrolling [34, 15, 19] reduce design complexity compared to traditional optimization solutions.
I-A3 Seamless Integration to 6G Networks
6G networks are expected to be AI-driven. In addition, sensing and localization capabilities will be integrated with traditional communication functions, creating intelligent networks that simultaneously map environments, track objects, and transmit data with higher efficiency [26]. Since ISAC is anticipated to play a key role in 6G systems, incorporating AI into ISAC designs is a natural step toward enabling its seamless adaptation to the evolving demands of 6G networks, making it intelligent by design.
The availability of real-world datasets is another important factor facilitating and motivating AI-based ISAC. Many ISAC applications, such as vehicle-to-everything systems [35, 36, 37], naturally generate large datasets, which enable data-driven learning and model fine-tuning. This, in turn, makes AI-based ISAC designs more practical and supports consistent benchmarking. Examples of such datasets include DeepSense 6G [38], a large-scale real-world dataset featuring synchronized communication, radar, camera, and GPS data collected in both indoor/outdoor scenarios; and the ISAC UAV Dataset [39], an outdoor radar-based channel measurements along with ground-truth drone trajectories, ideal for radar-centric ISAC tasks.
I-B Related Tutorials and Paper Scope
We review related survey and tutorial papers, outline the scope of our work, and highlight how it differs from existing surveys and tutorials.
I-B1 Existing Work
Several papers listed in Table I have surveyed prior research efforts contributing to the advancement of ISAC technology. The integration of radar/radio sensing into mobile networks is explored in [1], which surveys radar-centric, communication-centric, and joint design approaches. The fundamental limits of ISAC have been studied in [3] and [29]. In particular, [29] provides an in-depth discussion of key performance metrics, such as the CRLB and MSE, along with relevant theoretical limits, whereas [3] explores rate-distortion tradeoff as well as detection and estimation bounds.
All papers in Table I have topics related to signal/waveform design. However, [27] specifically focuses on this topic and surveys orthogonal frequency-division multiplexing (OFDM) and orthogonal time frequency space (OTFS) modulation for ISAC. OFDM-based sensing is also discussed in [26] and [29]. A recent tutorial on sensing with OFDM signals if offered in [22] which also includes a discussion on constellation design.
The scope of some of these papers extends beyond sensing and communication. Specifically, [26] not only reviews network sensing and sensing-aided communication techniques for cellular systems, but also addresses positioning, localization, and mapping. On the other hand, [23] also considers the role of computing, discussing integration gains, system and signal design, and network resource management for joint sensing, communication, and computation.
Finally, the scope of some other papers is more narrowly focused. For example, [25] presents a survey of interference management techniques for ISAC. In [28], intelligent multi-modal sensing–communication integration is explored using the concept of machine synesthesia. Meanwhile, [24] surveys secure ISAC for intelligent connectivity, emphasizing the importance of dynamic adaptability and self-learning capabilities, and [40] is about ISAC layered structure and standardization.
I-B2 Paper Scope
@@
This work presents a tutorial dedicated to AI-based ISAC, with a primary focus on how AI can enhance ISAC design through the optimization of waveform, constellation and beamforming. We discuss various ML/DL–based optimization approaches for theses design tasks, emphasizing their potential advantages over traditional analytical and numerical methods typically used for multi-objective optimization.
Existing ISAC surveys are typically centered around signal processing and multi-objective optimization methods. Although some have explored the use of DL in the ISAC context (see Table I), they primarily apply learning techniques to channel estimation. In contrast, this work aims to familiarize readers with a broader range of ML/DL-based approaches, with an emphasis on constellation, waveform, and beamforming design in ISAC systems.
@@ I-C Contribution
We begin with a primer on ISAC, introducing key concepts in both communication and sensing, along with performance metrics relevant to each domain. Radar technology is presented as a core sensing component, and its integration with communication systems is discussed. We then frame ISAC problems as multi-objective optimization tasks, comparing conventional approaches—such as weighted sum, $\epsilon$-constraint, and Pareto optimization—with modern learning-based solutions, including multi-task learning and multi-objective learning.
A major contribution of this tutorial is its detailed examination of waveform, beamforming, and constellation design for ISAC and the role AI plays in enhancing it. We cover traditional beamforming techniques in communication, radar, and ISAC and explore learning-based approaches, including supervised, unsupervised, and reinforcement learning, as well as model-based learning techniques like plug-and-play methods and algorithm unrolling.
To make AI-driven ISAC designs more tangible, we present three concrete case studies on AI-powered waveform, beamforming, and constellation design.
@@
  • The second case study presents a hybrid beamforming method using a learned optimizer, which optimizes analog and digital precoders by reformulating the problem as constrained least-squares optimization, and solves it via neural network-based projection. Numerical results show that the proposed unfolded hybrid beamforming method achieves spectral efficiency close to optimal digital beamforming, particularly with more receiving antennas.
  • The third case study employs a combination of end-to-end learning for communication and supervised learning for sensing. Specifically, an autoencoder-based architecture is designed and trained to learn a modulation constellation that achieves low bit error rates. Besides, a radar detector is trained using labeled data to identify the presence or absence of a target. The shared encoder enables a single transmitted signal to serve both functions, and allows balancing sensing and communication performance through joint optimization.
The three case studies collectively highlight AI’s adaptability across key ISAC physical-layer designs. The first applies unsupervised learning for waveform optimization, embedding sensing–communication trade-offs into the loss function. The second demonstrates a learned optimizer for hybrid beamforming, efficiently approximating constrained solutions. The third combines end-to-end and supervised learning for joint constellation design and radar detection, enabling unified signal generation. Together, they showcase AI’s potential for data-driven waveform, beamforming, and constellation design for ISAC in next-generation systems.
Finally, we elaborate on the lessons learned and outline potential future research directions. The Lessons Learned section includes detailed discussions on: (i) when to use each AI model; (ii) the benefits of AI-empowered ISAC systems; and (iii) the limitations of AI-based ISAC systems. The Future Research section covers promising topics such as the application of transformer neural networks for unified waveform and beamformer design, and digital twin-aided ISAC architectures. By bridging the gap between traditional signal processing and modern AI-driven techniques, this tutorial serves as a valuable resource for researchers and practitioners at the forefront of ISAC technology.
TABLE II: List of Key Abbreviations.
@@
PSK Phase-shift keying
PAPR Peak-to-average power ratio
PGA Projected gradient ascent
PnP Plug-and-play
@@
ReLU Rectified linear unit
RCS Radar cross section
@@ II A Primer to ISAC
TABLE III: Common Performance Metrics for Communication Systems
Metric Definition and Description References
BER/SER/BLER Probability of bit or symbol, block errors in received data. Lower values indicate more reliable communication. These metrics depend on ${E_{b}}/{N_{0}}$ (or SNR), modulation scheme, and coding. [41]
Spectral efficiency Bits transmitted per second per Hz of bandwidth. Higher spectral efficiency allows for more data transmission in limited bandwidth. [41]
Capacity Theoretical upper bound on data transmission rate. For a point-to-point channel, Shannon’s formula $\rm C=W\,\mathrm{log}_{2}\left(1+SNR\right)$ gives the maximum possible data rate with vanishing errors. [41]
Achievable rate Data transmission rate considering practical constraints. Determined by modulation, coding, and BER requirements. It is lower than capacity because of implementation limitations. [41]
Throughput Actual data successfully transmitted and received. Affected by network congestion, re-transmissions, and protocol overhead. It is always lower than the achievable data rate. [42, 43, 44]
Outage probability Probability that the received signal falls below a threshold. Important in fading channels to assess communication reliability. [45, 42]
Latency Delay in data transmission and reception. Low latency is crucial for real-time applications like voice, video, gaming, and autonomous systems. [46]
EVM Deviation between actual and ideal signal constellation points. Lower EVM indicates better modulation accuracy and signal quality. [47, 48, 49]